OpenStudy (anonymous):

If the unit price for watermelon is $0.24 per pound and Jared pays $1.68 for a watermelon, how many pounds did his watermelon weigh?

OpenStudy (anonymous):

6, 7, 8, or 4 pounds

OpenStudy (nurali):

1.68/.24 = 7

OpenStudy (anonymous):

Ok I will tell you if it is right, but it is a test so it might take time but can you help with the others?

OpenStudy (nurali):

ok i will try

OpenStudy (anonymous):

Cool Thanks

OpenStudy (nurali):

please click best response.

OpenStudy (anonymous):

A photocopy was made of Julie’s watercolor painting. The photocopy resized the painting so that it measured 4 inches wide by 7 inches long. If the actual painting is 16 inches wide, what is its length? 28, 9.14, 38, or 64

OpenStudy (anonymous):

Can you help with that?

OpenStudy (nurali):

4 in.wide to 7 in. long = 16 in. wide = x in. long 4/7 = 16/x 4x = (16)(7) 4x = 112 x = 28 in. long

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Ok I have a lot because it is a test so thanks a lot

OpenStudy (nurali):

pleasure

OpenStudy (anonymous):

What is 38% of 150? 5.7, 56.5, 47, or 57

OpenStudy (nurali):

38% of 150 = (.38)(150) = 57

OpenStudy (anonymous):

Thanks a lot

OpenStudy (nurali):

pleasue

OpenStudy (anonymous):

What is the amount of sales tax owed on a $450 laptop computer if the tax rate is 8%? $36.00 $31.50 $45.00 $3.60

OpenStudy (nurali):

450 x 8% = (450)(.08) = $ 36.00

OpenStudy (anonymous):

Thanks, Ok the next couple of problems have no A.B.C.D but if you can show your work just a bit like 5x4=20 and then do the next step if you can.

OpenStudy (anonymous):

Find the unit price for each option shown below. Round to the nearest cent when necessary. Indicate which option is the better buy. Option I: 10 candy bars for $6.75 Option II: 12 candy bars for $7.25

OpenStudy (anonymous):

This is what it says on my paper Remember to show all of the steps that you use to solve the problem. You can use the comments field to explain your work. Your teacher will review each step of your response to ensure you receive proper credit for your answer.

OpenStudy (nurali):

sorry no idea u ak other.

OpenStudy (anonymous):

Ok While I do that one can you do this one?...

OpenStudy (anonymous):

Calculate the final price you would pay to buy jacket that originally sells for $65, but is discounted by 40%; you also have to pay 5% sales tax. And again if you can show your work just a bit if you can.

OpenStudy (anonymous):

And again thank you that helped me a lot.

OpenStudy (nurali):

soory no idea

OpenStudy (anonymous):

Ok well if you want to know which ones are right wait a sec so I can finish it may take a bit.

OpenStudy (anonymous):

What about this? Natalie picked 135 berries in 15 minutes. If she continues picking at that rate, how long will it take her to pick 486 berries?

OpenStudy (anonymous):

It may take some time there are several of tease kind of problems so sorry

OpenStudy (anonymous):

I think they are all right though

OpenStudy (anonymous):

What is $26 + 5% Do you know?

OpenStudy (anonymous):

Are you even there anymore?

OpenStudy (anonymous):

Oh good you are thanks

OpenStudy (anonymous):

That is weird though? so $25 adding 5% changes to 1.3?

OpenStudy (anonymous):

I saw that on mine to and thought I was doing it wrong but strange

OpenStudy (nurali):

5% of 26$ is 1.3

OpenStudy (anonymous):

Oh I am dumb sorry.

OpenStudy (nurali):

so for adding 5% u add 26+1.3=27.3

OpenStudy (anonymous):

How did I not realize that

OpenStudy (anonymous):

Ok thanks

OpenStudy (nurali):

pleasure now i am leaving to continues my work ok bye

OpenStudy (anonymous):

Ok bye and thanks for the help